Consider the following balanced equation: 4ZnS + 6O2 4ZnO + 4SO2
Determine the mass of zinc oxide produced when 16.7 g of zinc sulfide is combined with
1.26 × 1023 molecules of O2.
1
Expert's answer
2019-11-26T06:40:14-0500

 4ZnS + 6O2 4ZnO + 4SO2 

moles of ZnS=16.7/97.3=0.712

This means also 0.712 moles of zinc sulphide will be produced because the reaction ratio of ZnS to ZnO is 1:1

To get masss of ZnO produced ;multiply the number of moles by rfm to get

81.4*0.712=57.96g